# Why do Machine Learning Notebooks Crash?

### v1

@@ 5.3 Comparison with existing work
Frequent bugs in notebooks
De Santana et al. [8] find that environment and configuration issues are the most common bugs in Jupyter notebooks, often discussed on platforms like Stack Overflow. However, our study shows a higher prevalence of coding errors (52%), including API misuse, data confusion, and implementation issues. This difference likely reflects data sources: while the analysis from De Santana et al. focuses on user-reported issues, our study examines error outputs directly from executed notebooks. Our approach captures runtime failures, highlighting coding-related errors, whereas environment issues are often resolved before execution and thus less visible in our dataset.
ML/Python bugs
Our study finds that ML crashes are more frequent than general Python crashes (Kaggle: 71%, GitHub: 59%), in contrast with findings by Zhang et al. [11]. This discrepancy arises partly from differing definitions: they define ML bugs at a system level, encompassing all bugs in ML components of a system, while we focus specifically on bugs in code interacting with ML libraries. Additionally, both studies analyze error messages, but theirs are derived from job failures where code is tested, whereas ours come from cell outputs in notebooks, capturing issues during early development. This suggests that our study more directly reflects initial challenges faced by developers.
ML pipeline
Our study finds that the data preparation stage is the most crash-prone (GitHub: 29%, Kaggle: 26%) in the ML pipeline, similar to Islam et al. [7]. While their study includes bugs related to poor performance and incorrect outputs, not just crashes, and does not detail the impact of bugs within each stage, both studies highlight that data preparation remains the most challenging stage.
Root cause identification
@@
Our study identifies new root causes of crashes in ML notebooks, particularly those related to notebook-specific issues like out-of-order execution and previous cell error. While out-of-order execution has been linked to reproducibility issues in prior work [2], it has not been explored as a crash root cause. Additionally, by analyzing crash information, context code, and cell outputs, we can distinguish crashes caused by data confusion from API misuse. Error messages and cell outputs reveal details about data shapes, types, and structures that reflect users’ understanding of their data, helping us identify crash origins. Examining cell outputs and documentation cells also enables us to identify intentional errors, such as API misuse shown in tutorial notebooks for instructional purposes.
@@ 6.1 External validity
Temporal limitation in GitHub dataset
The GitHub dataset includes notebooks from January 1, 2015, to March 31, 2022, which may not fully represent current notebook trends. Given the rapid evolution of coding practices and the tools used over time, notebooks from earlier years may reflect outdated coding styles or libraries. To mitigate this, we apply a filtering process to remove notebooks using older Python versions, improving the generalizability of the dataset.
Comparison of GitHub and Kaggle notebooks
The Kaggle dataset, which spans from January 1, 2023, to January 1, 2024, provides a more up-to-date snapshot of notebook trends. While the differing periods of GitHub and Kaggle datasets reduce redundancy in notebooks and associated crashes, the validity of comparison is threatened because ML practices may have changed. However, there is limited evidence of this effect because our results show no significant difference in exception types and root causes between Kaggle and GitHub.
@@ 6.2 Internal validity
Tutorial notebooks
The analyzed dataset contains various tutorial and educational notebooks. These tutorial notebooks often contain intentional crashes for didactic purposes. Such crashes could distort the crash patterns we aim to uncover. However, while these crashes are intentional, they still represent errors that users commonly believe occur in real-world scenarios, thus minimally affecting our results.
@@
Reproducibility of crashes
The error outputs we analyze represent real execution crashes but they are not necessarily reproducible due to factors like execution order, missing datasets, or environment replication [2]. While reproducing crashes could further validate our findings, it is challenging, especially for GitHub notebooks. Kaggle notebooks, with their accessible datasets and pre-configured environments, may enable reproducibility. In future work, we plan to reproduce crashes in Kaggle notebooks to validate our results and create a reproducible dataset for deeper analysis.
Clustering crashes
@@ 7 Conclusions and Future Work
In this paper, we present the first focused study on software crashes in ML notebooks from GitHub and Kaggle, uniquely identifying notebook-specific root causes, such as out-of-order execution and previous cell error. Our analysis classifies exception types, root causes, and ML pipeline stages, distinguishing between ML-specific and general Python bugs. Our findings reveal that the most common exceptions in ML notebooks are NameError, ValueError, and TypeError. The primary root causes are API misuse, data confusion, and notebook-specific issues (especially out-of-order execution). Notably, ML bugs account for more crashes than Python bugs, especially in Kaggle notebooks. Most crashes (58%) occur during the data preparation, model training, and evaluation/prediction stages of the ML pipeline. Additionally, GitHub and Kaggle exhibit different crash distributions across ML pipeline stages.
Future work will expand on these findings by constructing a reproducible dataset of ML crashes. Motivated by the results of this study, we also aim to develop an integrated, real-time bug detection tool tailored to ML development in the notebook environment, building on our previous work [39]. This tool will focus on ML bug detection while monitoring cell execution order and maintaining notebook semantics, helping developers catch errors as they code and enhancing productivity in ML development in notebooks.

### v2

@@ 5.3 Comparison with existing work
To contextualize our results, we compare our findings with previous studies [13, 10]. Zhang et al. [13] identify failure types for DL jobs similar to our crash types. They also report the distribution of failures across ML pipeline stages. De Santana et al. [10] study bugs in Jupyter notebooks at a general level, which are similar to our target data format. To enable a fair comparison, we map our results to existing classifications of failure types [13] and bug types [10]. This is unnecessary for ML pipeline stages as our categorization includes the categories used in the other study [13].
For the study by Zhang et al. [13], we retain only DL notebooks to match their domain. This filtering process follows a similar approach to identifying ML notebooks, as described in Section 3.2, where we compile a list of DL libraries as shown in Table II. Following the same methodology as ML library identification, two authors independently label the DL libraries, achieving a 100% agreement rate. The overall mapped results are presented in Table VII, while details on our mapping strategy and subcategory comparisons are available in our GitHub repository [15].
Bug type comparison with Zhang et al. [13]
As VII(a) shows, execution environment errors are substantially lower in our DL dataset (compared to Zhang et al. [13]). The authors attribute the high frequency of execution environment errors to differences between job execution environments and engineers’ local environments. Our findings indirectly support this claim. Our target artifacts are DL notebooks typically executed either locally (GitHub) or in controlled environments with pre-installed dependencies (Kaggle). When developing DL notebooks in such stable environments, environment-related bugs are less common. Transitioning to other environments significantly increases the likelihood of encountering environment errors.
Takeaway 7: Crashes caused by execution environment are less frequent in ML notebooks but they can emerge when transitioning to execution platforms (e.g., cloud-based) used for deep learning. Practitioners should proactively test for such issues to ensure smooth deployment.
After excluding the effect of execution environment errors, our study still identifies a higher prevalence of DL-specific crashes. This discrepancy can be attributed to the different development stages being analyzed. Our results derive from examining cell outputs in notebooks primarily used for prototyping and early-stage DL development. In contrast, Zhang et al. mainly focus on later development stages when jobs (tested locally) are already being submitted to a DL platform. As such, our study should better represent the initial difficulties faced by DL developers.
Takeaway 8: As ML notebooks are used in early stage of development, deep learning-specific crashes are more prevalent in ML notebooks compared to deep learning jobs.
Bug type comparison with De Santana et al. [10]
As shown in VII(b), De Santana et al. [10] find that environment and settings errors are considerably more common in Jupyter notebooks, as observed in GitHub commits, compared to our findings. This discrepancy is likely due to differences in the nature of the analyzed notebooks. Specifically, their dataset primarily consists of notebooks from stable and mature GitHub projects, such as AWS, Azure, Fastai, and TensorFlow, whereas our dataset focuses on notebooks used for ML prototyping and experiments. To better understand this difference, we randomly sample 20 commits from their dataset to observe that the associated notebooks are predominantly demonstrations, documentation, or usage examples of specific libraries or software created by the developers of those libraries. The corresponding commits are mostly aimed at updating dependencies required by the projects rather than fixing implementation bugs, reflecting the stable and mature nature of these projects. As a result, it is expected that environment and settings issues are more prevalent in their dataset.
@@
We believe that our study better reflects the practical development challenges in notebook environments. We observe unique implementation errors, particularly those related to notebook-specific issues like out-of-order execution and previous cell errors. While out-of-order execution has been linked to reproducibility concerns in prior research [2], its role as a root cause of crashes has not been widely explored. By analyzing crash information, code, and cell outputs, our study also distinguishes crashes caused by data confusion from those resulting from API misuse. Error messages and outputs provide crucial insights into data shapes and structures, helping us pinpoint the origin of crashes more effectively. Therefore, we present these implementation issues separately in VII(b) by marking them as “Implementation (other).”
(a) Bug type classification comparison with Zhang et al. (S1 [13]). The last column shows normalized percentages excluding execution environment (Env.). Our results are based on DL-only notebooks crashes.
Dimensions S1 [13] Result S1 [13] Result Ours (DL) Exclude Env. S1 [13] vs. Ours
Execution environment 48.0% 19.9% -
Data 2.0% 0.9% 3.8% vs. 1.1%
DL specific 13.5% 32.5% 26.0% vs. 40.6%
General code error 36.5% 46.7% 70.2% vs. 58.3%
(b) Bug type classification comparison with De Santana et al. (S2 [10]). We present our mapped results separately for Kaggle and GitHub for fair comparison.
Dimensions S2 [10] Result S2 [10] (GH) Result Ours (KG) Result Ours (GH)
Environment and settings 35.6% 3.4% 6.9%
Processing 1.9% 3.1% 1.5%
Implementation 44.2% 26.4% 23.1%
Implementation (other) - 67.1% 68.5%
Other (e.g., Kernel) 18.3% - -
(c) ML pipeline stage comparison with S1 [13] (DL cases only). Our normalized percentages exclude data visualization (DV) and model construction (MC) stages absent in S1.
ML pipeline stages Our category Result S1 [13] Exclude DV&MC Ours (DL) Result Ours (DL)
Environment setup 30.8% 10.0% 7.6%
Data preparation 30.0% 22.7% 17.3%
Data visualization - - 7.6%
Model construction - - 16.4%
Training 15.0% 35.7% 27.2%
Evaluation/prediction 24.3% 31.6% 24.0%
TABLE VII: Comparison with existing studies [13, 10]. An overview of these studies is provided in Table I.
Takeaway 9: The error patterns prevalent in Jupyter notebooks created by authors of ML libraries are different from ML notebooks created by the users of such libraries.
ML pipeline comparison with Zhang et al. [13]
As VII(c) shows, Zhang et al. [13] report considerably more environment setup failures but fewer training failures than our results. Their higher environment setup failure rate aligns with their larger proportion of execution environment bugs (VII(a)). Fewer training errors may reflect that DL developers focus more on ensuring training correctness when developing locally before job submission.
As a novel finding, our analysis reveals that data visualization can also be a source of crashes in ML workflows because notebooks serve not only for coding but also for documentation and exploratory data analysis. The omission of this stage in prior work suggests that visualization-related crashes might be underrepresented. Identifying this category highlights the unique nature of notebook development and underscores the need for debugging support tailored to interactive workflows.
Takeaway 10: Since ML notebooks blend coding with exploration and documentation, notebook development has some unique types of errors uncommon in traditional ML workflows, particularly, in data visualization.
@@ 6.1 External validity
Data selection bias
As our dataset only includes publicly shared GitHub and Kaggle notebooks, it may exclude crashes that users fix before publishing, thus crashes encountered during development may be underrepresented. As a result, our findings reflect crashes that persist, which may not be the full spectrum of debugging challenges. Future work could address this by analyzing version histories or autosaved checkpoints to capture a more complete development process.
Temporal limitations in datasets
The GitHub dataset includes notebooks from January 1, 2015, to March 31, 2022, while the Kaggle dataset spans from January 1, 2023, to January 1, 2024. Given the rapid evolution of coding practices and tools used over time, notebooks from earlier years may reflect outdated coding styles or libraries. To mitigate this, our filtering process removes notebooks using older Python versions, improving the generalizability of the dataset.
@@ 6.2 Internal validity
Clustering crashes
Sampling limitations
Our proportional stratified sampling ensures broad representation, but rare strata may still be underrepresented despite a minimum sample size of one per stratum. Increasing this minimum was impractical due to manual labeling constraints. To assess potential gaps, we conduct an exploratory analysis by manually examining a maximum of 10 crashes from 10 randomly selected small clusters in GitHub and Kaggle datasets separately. This reveals no previously unknown crash patterns, suggesting that our approach captures key crash types and root causes, although some rare errors may still be missed. Furthermore, cases excluded during resampling reduce certain strata sizes, which may impact the original sampling proportions.
@@
Reproducibility of crashes and root cause analysis
The error outputs we analyze represent real execution crashes, but they are not necessarily reproducible due to factors like execution order, missing datasets, or environment replication [2]. While reproducing crashes could further validate our findings, it is challenging, especially for GitHub notebooks. Kaggle notebooks, with their accessible datasets and pre-configured environments, may enable reproducibility. Nevertheless, we carried out a small-scale validation to assess the accuracy of our root cause analysis by trying to reproduce a few crashes in Kaggle, and the high accuracy suggests that our approach is reliable (see Section 4.1). In future work, we plan to extend crash reproducibility efforts in Kaggle notebooks to validate our results thoroughly and create a reproducible dataset for deeper analysis.
@@ 7 Conclusions and Future Work
This paper provides the first large-scale empirical study focusing on software crashes in public ML notebooks collected from GitHub and Kaggle. Our analysis classifies (1) crash types, (2) root causes, (3) ML pipeline stages, and (4) involved ML libraries. (5) We also distinguish between ML-specific and general Python bugs.
Our findings reveal that the (1) most common crashes in ML notebooks are Variable Not Found, Invalid Argument, and IO Error. We have also identified prevalent ML-specific crash types such as Tensor Shape Mismatch, Data Value Violation, and Out of Memory. Moreover, (2) the primary root causes of crashes are API misuse and notebook-specific issues (specifically, out-of-order execution and previous cell error). (3) Most crashes (70.8%) occur during data preparation, model training, and evaluation/prediction stages of the ML pipeline, while data visualization bugs are observed for the first time. (4) We also identify that certain ML libraries are more challenging to use for ML notebook development, in fact, crashes mostly occur when using DL libraries tensorflow/keras and torch. Finally, (5) ML bugs account for more crashes than Python bugs, especially in Kaggle notebooks. Future work will expand on these findings by constructing a reproducible dataset of ML crashes, and investigate the temporal evolution of crash patterns.